1/ Given the following data: sales $1,500,000; gross profit $640,000; net income after tax $40,000 and income tax expense $35,000. What is the common-size percentage for operating expenses?

37.7%

42.7%

95.0%

97.3%

2/ To best interpret the accounts receivable turnover ratio, the days in accounts receivable should be compared to the company's

inventory turnover.

sales revenue.

credit terms.

accounts receivable balance.

3/ The quick ratio will be negatively impacted by

tying up cash in inventory.

increasing accounts receivable.

decreasing the level of prepaid accounts.

increasing levels of long term debt.

4/ Which of the following depicts the current ratio?

(current assets – inventory) ÷ current liabilities

currents assets ÷ total assets

(current assets – inventory) ÷ total assets

current assets ÷ current liabilities
Answer #1

1.

Net income after tax = $40,000

Income tax expense = $35,000

Income before tax = Net income after tax + Income tax expense

= 40,000+35,000

= $75,000

Operating expenses = Gross profit - Income before tax

= 640,000-75,000

= $565,000

Common size percentage for operating expenses = Operating expenses/sales

= 565,000/1,500,000

= 37.7%

First option is correct option.

2. To best interpret the accounts receivable turnover ratio, the days in accounts receivable should be compared to the company's credit terms.

Third option is correct option.

3.

The quick ratio will be negatively impacted by tying up cash in inventory.

First option is correct option.

When more and more of inventory is purchased on cash, quick assets decrease due to decrease in cash and thus quick ratio declines.

4.

Current ratio = Current assets/Current liabilities

Fourth option is correct option.

Kindly comment if you need further assistance.

Thanks‼!